# Prox-DBRO-VR: A Unified Analysis on Decentralized Byzantine-Resilient Composite Stochastic Optimization with Variance Reduction and Non-Asymptotic Convergence Rates

### v6

@@ III-B A General Algorithmic Framework
The Byzantine resilience of Prox-DBRO-VR is attained by adopting the robust consensus aggregation based on total variation, which is initially studied in [45]. The literature [28, 33] extends this strategy to handling distributed federated learning problems, and [27] studies it in a decentralized manner. However, all these works [28, 27, 33] not only rely on a bounded-variance assumption in theoretical analysis, but establish slower sub-linear convergence rates. Thus, the most important goal of designing Prox-DBRO-VR is to achieve linear convergence and removes the bounded-variance assumption, which can be attained with the aid of VR techniques.
@@ III-C Prox-DBRO-SAGA and Prox-DBRO-LSVRG
We introduce the localized version of two popular centralized VR techniques, SAGA [41] and LSVRG [42], into Prox-DBRO-VR, to develop Prox-DBRO-SAGA and Prox-DBRO-LSVRG. The detailed updates of Prox-DBRO-SAGA and Prox-DBRO-LSVRG are presented in Algorithms 2-3, respectively.
Algorithm 2 Prox-DBRO-SAGA
0:  Each reliable agent $i$, $i\in\mathcal{R}$, initializes the same parameters and starting points according to Algorithm 1, and auxiliary variables $u_{i,1}^{l}=u_{i,0}^{l}={x_{i,0}},\forall l\in\mathcal{Q}_{i}$, together with gradient tables $\left\{{\nabla f_{i}^{l}\left({u_{i,0}^{l}}\right)}\right\}_{l=1}^{{q_{i}}}$ additionally.
1:  for all $k=0,1,2,\ldots$ do
2:     Each reliable agent $i$, $i\in\mathcal{R}$, exchanges information according to Step 2 in Algorithm 1.
3:     Each reliable agent $i$, $i\in\mathcal{R}$, selects uniformly a random sample with index $s_{i,k}$ from the set $\mathcal{Q}_{i}$ and evaluates the lcoal stochastic gradient
@@
4:     Each reliable agent $i$, $i\in\mathcal{R}$, takes $u_{i,k+1}^{s_{i,k}}={x_{i,k}}$ and replaces $\nabla f_{i}^{{s_{i,k}}}\left({u_{i,k+1}^{s_{i,k}}}\right)$ by $\nabla f_{i}^{{s_{i,k}}}\left({{x_{i,k}}}\right)$ in the corresponding position of the gradient table, while keeps $\nabla f_{i}^{l}\left({u_{i,k+1}^{l}}\right)=\nabla f_{i}^{l}\left({u_{i,k}^{l}}\right)$, $l\in{\mathcal{Q}_{i}}\backslash\left\{{{s_{i,k}}}\right\}$.
5:     Each reliable agent $i$, $i\in\mathcal{R}$, updates its current model according to Steps 4-5 in Algorithm 1.
6:  end for
Algorithm 3 Prox-DBRO-LSVRG
0:  Each reliable agent $i$, $i\in\mathcal{R}$, initializes the same parameters and starting points according to Algorithm 1, and an auxiliary variable ${w_{i,0}}={x_{i,0}}$ additionally.
1:  for all $k=0,1,2,\ldots$ do
2:     Each reliable agent $i$, $i\in\mathcal{R}$, exchanges information according to Step 2 in Algorithm 1.
3:     Each reliable agent $i$, $i\in\mathcal{R}$, selects uniformly a random sample with index $s_{i,k}$ from the set $\mathcal{Q}_{i}$ and evaluates the local stochastic gradient
@@
4:     Each reliable agent $i$, $i\in\mathcal{R}$, takes $w_{i,k+1}={x_{i,k}}$ with an uncoordinated triggered probability $p_{i}$ and keeps $w_{i,k+1}={w_{i,k}}$ with the probability $1-p_{i}$.
5:     Each reliable agent $i$, $i\in\mathcal{R}$, updates its current model according to Steps 4-5 of Algorithm 1.
6:  end for
@@
Note that all steps in Prox-DBRO-SAGA and Prox-DBRO-LSVRG, together with Prox-DBRO-VR, are executed in parallel among all reliable agents. It is also worthwhile to mention that the expected cost in evaluating the stochastic gradient under Prox-DBRO-LSVRG is at least double that of Prox-DBRO-SAGA at every iteration. However, this computational advantage of Prox-DBRO-SAGA is at the expense of an expensive storage cost of ${\cal O}\left({n{q_{i}}}\right)$ for each agent $i$ owing to the employment of the gradient table, while Prox-DBRO-LSVRG does not incur extra storage to save the local batch gradients. Therefore, adopting either Prox-DBRO-SAGA or Prox-DBRO-LSVRG in practice involves a trade-off between per-iteration computational cost and storage. Users can also implement Prox-DBRO-VR via incorporating other categories of VR techniques based on their customized needs.
@@ IV Convergence Analysis
For the simplicity of notation, we denote ${{\mathcal{F}_{k}}}$ as the filter of the history with respect to the dynamical system generated by the sequence $\left\{{s_{k}^{i}}\right\}_{k\geq 0}^{i=1,2,\ldots,\left|\mathcal{R}\right|}$, and the conditional expectation $\mathbb{E}\left[{{s_{k}}|{\mathcal{F}_{k}}}\right]$ is shortly denoted by ${\mathbb{E}_{k}}\left[\cdot\right]$ in the sequel analysis. To facilitate the subsequent analysis, we give the sequel definitions.
${x_{k}}:=$ ${\left[{x_{1,k}^{\top},\ldots,x_{\left|\mathcal{R}\right|,k}^{\top}}\right]^{\top}}\in{\mathbb{R}^{\left|\mathcal{R}\right|n}},$
${r_{k}}:=$ ${\left[{r_{1,k}^{\top},\ldots,r_{\left|\mathcal{R}\right|,k}^{\top}}\right]^{\top}}\in{\mathbb{R}^{\left|\mathcal{R}\right|n}},$
$\nabla F\left({{x_{k}}}\right):=$ ${\left[{\nabla{f_{1}}{{\left({{x_{1,k}}}\right)}^{\top}},\ldots,\nabla{f_{\left|\mathcal{R}\right|}}{{\left({{x_{\left|\mathcal{R}\right|,k}}}\right)}^{\top}}}\right]^{\top}}\in{\mathbb{R}^{\left|\mathcal{R}\right|n}},$
${\partial_{x}}G\left({{x_{k}}}\right):=$ $\left[{\partial_{{x_{1}}}}g{{\left({{x_{1,k}}}\right)}^{\top}},\ldots,{{\partial_{{x_{\left|\mathcal{R}\right|}}}}g{{\left({{x_{\left|\mathcal{R}\right|,k}}}\right)}^{\top}}}\right]^{\top}\!\in\!{\mathbb{R}^{\left|\mathcal{R}\right|n}},$
$\chi\left({{x_{k}}}\right):=$ $\sum\limits_{i\in\mathcal{R}}{{\phi_{i}}\sum\limits_{j\in{\mathcal{R}_{i}}}{{{\left\|{{x_{i,k}}-{x_{j,k}}}\right\|}_{a}}}}\in{\mathbb{R}},$
$\delta\left({{x_{k}}}\right):=$ $\sum\limits_{i\in\mathcal{R}}{{\phi_{i}}\sum\limits_{j\in{\mathcal{B}_{i}}}{{{\left\|{{x_{i,k}}-{z_{ij,k}}}\right\|}_{a}}}}\in{\mathbb{R}},$
${\partial_{{x_{i}}}}\chi\left({{x_{i,k}}}\right):=$ ${\phi_{i}}\sum\limits_{j\in{\mathcal{R}_{i}}}{{\partial_{{x_{i}}}}{{\left\|{{x_{i,k}}-{x_{j,k}}}\right\|}_{a}}}\in{\mathbb{R}^{n}},$
${\partial_{x}}\chi\left({{x_{k}}}\right):=$ ${\left[{{\partial_{{x_{1}}}}\chi{{\left({{x_{1,k}}}\right)}^{\top}},\ldots{\partial_{{x_{\left|\mathcal{R}\right|}}}}\chi{{\left({{x_{\left|\mathcal{R}\right|,k}}}\right)}^{\top}}}\right]^{\top}}\!\in\!{\mathbb{R}^{\left|\mathcal{R}\right|n}},$
${\partial_{{x_{i}}}}\delta\left({{x_{i,k}}}\right):=$ ${\phi_{i}}\sum\limits_{j\in{\mathcal{R}_{i}}}{{\partial_{{x_{i}}}}{{\left\|{{x_{i,k}}-{z_{ij,k}}}\right\|}_{a}}}\in{\mathbb{R}^{n}},$
${\partial_{x}}\delta\left({{x_{k}}}\right):=$ ${\left[{{\partial_{{x_{1}}}}\delta{{\left({{x_{1,k}}}\right)}^{\top}},\ldots,{\partial_{{x_{\left|\mathcal{R}\right|}}}}\delta{{\left({{x_{\left|\mathcal{R}\right|,k}}}\right)}^{\top}}}\right]^{\top}}\!\in\!{\mathbb{R}^{\left|\mathcal{R}\right|n}}.$
Based on these definitions, we next provide briefly a compact form of Prox-DBRO-VR for the subsequent convergence analysis as follows:
@@ IV-A Auxiliary Results
Note that both sequences ${\left\{{t_{i,k}^{u}}\right\}_{i\in\mathcal{R},k\geq 0}}$ and ${\left\{{t_{i,k}^{w}}\right\}_{i\in\mathcal{R},k\geq 0}}$ are non-negative according to the convexity of the local component function ${f_{i}^{l}}$, $l\in\mathcal{Q}_{i}$. For the sequel analysis, we define respectively the gradient-learning quantities $t_{k}^{u}:=\sum\nolimits_{i\in\mathcal{R}}{t_{i,k}^{u}}$ and $t_{k}^{w}:=\sum\nolimits_{i\in\mathcal{R}}{t_{i,k}^{w}}$ for Prox-DBRO-SAGA and Prox-DBRO-LSVRG, the largest and smallest number of local samples ${q_{\min}}:={\min_{i\in\mathcal{R}}}{q_{i}}$ and ${q_{\max}}:={\max_{i\in\mathcal{R}}}{q_{i}}$, the minimum and maximum triggered probabilities ${p_{\min}}:={\min_{i\in\mathcal{R}}}{p_{i}}$ and ${p_{\max}}:={\max_{i\in\mathcal{R}}}{p_{i}}$, while ${\kappa_{q}}:={q_{\max}}/{q_{\min}}\geq 1$.
@@
(Gradient-Learning Quantity) Suppose that Assumptions 1-2 hold. For $k\geq 0$, we have for Prox-DBRO-SAGA,
@@
where ${D_{F}}\left({{x_{k}},{x^{*}}}\right):=F\left({{x_{k}}}\right)-F\left({{x^{*}}}\right)-\nabla F{\left({{x^{*}}}\right)^{\top}}\left({{x_{k}}-{x^{*}}}\right)$ is known as the Bregman divergence of the convex cost function $F$ due to the convexity preservation.
@@
We next seek the upper bound of the distance between the stochastic gradient estimator $r_{k}$ and the optimal gradient $\nabla F\left({{x^{*}}}\right)$ for both Prox-DBRO-SAGA and Prox-DBRO-LSVRG.
@@
(Gradient-Learning Error) Suppose that Assumptions 1-2 hold. For $k\geq 0$, we have for Prox-DBRO-SAGA,
@@
where $r_{k}^{u}:={\left[{{{\left({r_{1,k}^{u}}\right)}^{\top}},{{\left({r_{2,k}^{u}}\right)}^{\top}},\ldots,{{\left({r_{\left|\mathcal{R}\right|,k}^{u}}\right)}^{\top}}}\right]^{\top}}\in{\mathbb{R}^{\left|\mathcal{R}\right|n}}$ and $r_{k}^{w}:={\left[{{{\left({r_{1,k}^{w}}\right)}^{\top}},{{\left({r_{2,k}^{w}}\right)}^{\top}},\ldots,{{\left({r_{\left|\mathcal{R}\right|,k}^{w}}\right)}^{\top}}}\right]^{\top}}\in{\mathbb{R}^{\left|\mathcal{R}\right|n}}$.
@@
Consider two positive constants $a\geq 1$ and $b$, such that $11/{a}+1/{b}=$ 1. For an arbitrary vector $\tilde{x}\in{\mathbb{R}^{n}}$, we denote the sub-differential ${\partial}{\left\|{\tilde{x}}\right\|_{a}}=\left\{{\tilde{z}\in{\mathbb{R}^{n}}:\left\langle{\tilde{z},\tilde{x}}\right\rangle={{\left\|{\tilde{x}}\right\|}_{a}},{{\left\|{\tilde{z}}\right\|}_{b}}\leq 1}\right\}$.
@@
where $x=\left[{x_{1}^{\top},x_{2}^{\top},\ldots,x_{\left|\mathcal{R}\right|}^{\top}}\right]^{\top}\in{\mathbb{R}^{\left|\mathcal{R}\right|n}}$ and $y=\left[{y_{1}^{\top},y_{2}^{\top},\ldots,y_{\left|\mathcal{R}\right|}^{\top}}\right]^{\top}\in{\mathbb{R}^{\left|\mathcal{R}\right|n}}$.
@@ IV-B Main Results
We next derive a feasible range for the uncoordinated penalty parameters to enable the equivalence between the decentralized consensus optimization problem (3) and norm-penalized approximation problem (4) as follows, which further guarantees the equivalence between the original optimization problem (1) and norm-penalized approximation problem (4).
@@
(Robust Consensus Condition) Suppose that Assumptions 1 and 2 hold. Given any $g_{\tilde{x}}^{*}\in{\partial_{\tilde{x}}}{g}\left({{{\tilde{x}}^{*}}}\right)$, if we choose $\underline{\phi}\geq{\phi_{\min}}:={\left|\mathcal{R}\right|^{\frac{3}{2}}}\sqrt{\left|{{\mathcal{E}_{\mathcal{R}}}}\right|}\mathop{\max}\nolimits_{i\in\mathcal{R}}{\left\|{\nabla{f_{i}}\left({{{\tilde{x}}^{*}}}\right)}+{g_{\tilde{x}}^{*}}\right\|_{\infty}}/{\lambda_{\min}}\left(\Pi\right)$, the optimal solution to the original optimization problem (1) is equivalent to the globally optimal solution to norm-penalized approximation problem (4), i.e., ${x^{*}}={1_{\left|\mathcal{R}\right|}}\otimes{{\tilde{x}}^{*}}$.
@@
Theorem 1 demonstrates that a selection of sufficiently-large uncoordinated penalty parameters guarantees the equivalence between the original optimization problem (1) and norm-penalized approximation problem (4). However, the sequel convergence results manifest that a larger $\underline{\phi}$ causes a bigger convergence error. Therefore, the notion of sufficiently large uncoordinated penalty parameters is tailored for theoretical results, and one can hand-tune this parameter to obtain better algorithm performances in practice.
For simplicity, we fix the minimum and maximum uncoordinated triggered probabilities as ${p_{\min}}=1/{q_{\max}}$ and $p_{\max}=1/{q_{\min}}$, respectively. The following analysis considers ${r_{k}}:=\left\{\begin{gathered}r_{k}^{u},{\text{for {Prox-DBRO-SAGA}}}\hfill\\ r_{k}^{w},{\text{for {{Prox-DBRO-LSVRG}}}}\hfill\\ \end{gathered}\right.$ and ${t_{k}}:=\left\{\begin{gathered}t_{k}^{u},\text{for {Prox-DBRO-SAGA}}\hfill\\ t_{k}^{w},\text{for {Prox-DBRO-LSVRG}}\hfill\\ \end{gathered}\right.$, such that the theoretical results for both Prox-DBRO-SAGA (Algorithm 2) and Prox-DBRO-LSVRG (Algorithm 3) can be unified in a general framework. Before deriving a linear convergence rate for Algorithms 2-3, we first define the sequel parameters: $\underline{\phi}:={\min_{i\in\mathcal{R}}}{\phi_{i}}$, $\bar{\phi}:={\max_{i\in\mathcal{R}}}{\phi_{i}}$, $\gamma:=\mu L/\left({\mu+L}\right)$, $P_{1}^{c}:=16n{\bar{\phi}^{2}}\sum\nolimits_{i\in{\cal R}}{{{\left|{{{\cal R}_{i}}}\right|}^{2}}}+4n{\bar{\phi}^{2}}\sum\nolimits_{i\in{\cal R}}{{{\left|{{{\cal B}_{i}}}\right|}^{2}}}$, $P_{2}:=n{\bar{\phi}^{2}}\sum\nolimits_{i\in{\cal R}}{{{\left|{{{\cal B}_{i}}}\right|}^{2}}}/\gamma$, and $E:=4{P_{2}}/\gamma$.
@@
[…]
$\mathbb{E}\left[{\left\|{{x_{k}}-{1_{m}}\otimes{{\tilde{x}}^{*}}}\right\|_{2}^{2}}\right]$ (13)
@@
We continue to derive the sub-linear convergence rate of Algorithms 2-3 with the aid of the following bounded-gradient assumption on the non-smooth objective function $g$, which is standard in literature [49, 50].

### v12

@@ 6             Update its current local model according to the local proximal gradient step:
Inspired by [36], the Byzantine resilience of Prox-DBRO-VR is achieved by employing the resilient aggregation based on norm-penalized approximation. The literature [28, 34] leverages this strategy to handling distributed federated learning problems, and [27] studies it in a decentralized manner. However, all these algorithms [28, 27, 34] not only rely on a bounded-variance assumption in theoretical analysis but also exhibit a fundamental trade-off between convergence accuracy and speed. Therefore, one important motivation behind the design of Prox-DBRO-VR is to bypass the bounded-variance assumption while resolving the fundamental trade-off, which can be achieved with the aid of VR techniques.
@@ III-C Prox-DBRO-SAGA and Prox-DBRO-LSVRG
We introduce the localized version of two popular centralized VR techniques SAGA [43] and LSVRG [44], into Prox-DBRO-VR, to develop Prox-DBRO-SAGA and Prox-DBRO-LSVRG. The detailed updates of Prox-DBRO-SAGA and Prox-DBRO-LSVRG are presented in Algorithms 2-3, respectively.
Input: a proper constant or decaying step-size $\alpha_{k}>0$.
Initialize: the same parameters and starting points according to Algorithm 1 and auxiliary variables $u_{i,1}^{l}=u_{i,0}^{l}={x_{i,0}},\forall l\in\mathcal{Q}_{i}$, together with a gradient table $\left\{{\nabla f_{i}^{l}\left({u_{i,0}^{l}}\right)}\right\}_{l=1}^{{q_{i}}}$.
1 for $k=0,1,\ldots,$ do
2       for each reliable agent $i\in\mathcal{R}$ do
3             Exchange information according to Step 3 in Algorithm 1.
4             Select uniformly a random sample with index $s_{i,k}$ from the set $\mathcal{Q}_{i}$ and evaluate the local stochastic gradient according to
@@
5             Take $u_{i,k+1}^{s_{i,k}}={x_{i,k}}$ and replace $\nabla f_{i}^{{s_{i,k}}}\left({u_{i,k+1}^{s_{i,k}}}\right)$ by $\nabla f_{i}^{{s_{i,k}}}\left({{x_{i,k}}}\right)$ in the corresponding position of the gradient table, while keep the rest of the positions unchanged, i.e., $\nabla f_{i}^{l}\left({u_{i,k+1}^{l}}\right)=\nabla f_{i}^{l}\left({u_{i,k}^{l}}\right)$, $l\in{\mathcal{Q}_{i}}\backslash\left\{{{s_{i,k}}}\right\}$;
6             Update its current model according to Steps 7-8 in Algorithm 1.
7      for each Byzantine agent $i\in\mathcal{B}$ do
8            Send an arbitrary vector ${z_{ij,k}}$ to its neighbor $j$, $j\in{\mathcal{N}_{i}}$.
9
Output: all decision variables ${x_{i,k}}$, $i\in\mathcal{V}$, until a prescribed criterion is satisfied.
Algorithm 2 Prox-DBRO-SAGA.
Input: a proper constant or decaying step-size $\alpha_{k}>0$.
Initialize: the same parameters and starting points according to Algorithm 1 and an auxiliary variable ${w_{i,0}}={x_{i,0}}$.
1 for $k=0,1,\ldots,$ do
2       for each reliable agent $i\in\mathcal{R}$ do
3             Exchange information according to Step 3 in Algorithm 1.
4             Select uniformly a random sample with index $s_{i,k}$ from the set $\mathcal{Q}_{i}$ and evaluate the local stochastic gradient according to
@@
5             Take $w_{i,k+1}={x_{i,k}}$ with a heterogenous triggering probability $p_{i}$ and keep $w_{i,k+1}={w_{i,k}}$ with the probability $1-p_{i}$;
6             Update its current model according to Steps 7-8 in Algorithm 1.
7      for each Byzantine agent $i\in\mathcal{B}$ do
8            Send an arbitrary vector ${z_{ij,k}}$ to its neighbor $j$, $j\in{\mathcal{N}_{i}}$.
9
Output: all decision variables ${x_{i,k}}$, $i\in\mathcal{V}$, until a prescribed criterion is satisfied.
Algorithm 3 Prox-DBRO-LSVRG.
@@
All steps in Algorithms 1-3 are executed in parallel among all reliable agents since they are honest and hence comply with these update rules. It is also worthwhile to mention that the expected cost in evaluating the local stochastic gradient under Prox-DBRO-LSVRG is at least double that of Prox-DBRO-SAGA at every iteration. This computational advantage of Prox-DBRO-SAGA is at the expense of a higher (total) storage cost of ${\mathcal{O}}\left({nQ}\right)$ (with $Q=\sum\nolimits_{i=1}^{m}{{q_{i}}}$) than ${\mathcal{O}}\left(mn\right)$ of Prox-DBRO-LSVRG, which is prohibitively expansive when the number ($q_{i}$) of local training samples is massive. Therefore, adopting either Prox-DBRO-SAGA or Prox-DBRO-LSVRG in practice involves a trade-off between per-iteration computational cost and storage. Users may improve and implement Prox-DBRO-VR via incorporating other categories of VR techniques [49] based on their customized needs.
@@ IV Convergence Analysis
To streamline notation, we denote ${{\mathcal{F}_{k}}}$ as the filter of the history w.r.t. the dynamical system generated by the sequence ${\left\{{{s_{i,k}}}\right\}_{i\in\mathcal{R},k\geq 0}}$, and the conditional expectation $\mathbb{E}\left[{{s_{k}}|{\mathcal{F}_{k}}}\right]$ is shortly denoted by ${\mathbb{E}_{k}}\left[\cdot\right]$ in the sequel analysis. Let ${x_{k}}:={\mathbf{col}}{\left\{{{x_{i,k}}}\right\}_{i\in\mathcal{R}}}\in{\mathbb{R}^{\left|\mathcal{R}\right|n}}$ and ${r_{k}}:={\mathbf{col}}{\left\{{{r_{i,k}}}\right\}_{i\in\mathcal{R}}}\in{\mathbb{R}^{\left|\mathcal{R}\right|n}}$. To facilitate the subsequent analysis, we define
$\nabla F\left({{x_{k}}}\right):={\mathbf{col}}{\left\{{\nabla{f_{i}}\left({{x_{i,k}}}\right)}\right\}_{i\in\mathcal{R}}},$
${\chi_{i}}\left({{x_{i,k}}}\right):=\phi\sum\limits_{j\in{\mathcal{R}_{i}}}{{{\left\|{{x_{i,k}}-{x_{j,k}}}\right\|}_{a}}},\forall i\in\mathcal{R},$
${\partial_{x}}\chi\left({{x_{k}}}\right):={\mathbf{col}}{\left\{{{\partial_{{x_{i}}}}{\chi_{i}}\left({{x_{i,k}}}\right)}\right\}_{i\in\mathcal{R}}},$
${\delta_{i}}\left({{x_{i,k}}}\right):=\phi\sum\limits_{j\in{\mathcal{B}_{i}}}{{{\left\|{{x_{i,k}}-{x_{j,k}}}\right\|}_{a}}},\forall i\in\mathcal{R},$
${\partial_{x}}\delta\left({{x_{k}}}\right):={\mathbf{col}}{\left\{{{\partial_{{x_{i}}}}{\delta_{i}}\left({{x_{i,k}}}\right)}\right\}_{i\in\mathcal{R}}},$
such that Prox-DBRO-VR can be concisely formulated as:
@@ IV-A Auxiliary Results
Note that both sequences ${\left\{{t_{i,k}^{u}}\right\}_{i\in\mathcal{R},k\geq 0}}$ and ${\left\{{t_{i,k}^{w}}\right\}_{i\in\mathcal{R},k\geq 0}}$ are non-negative according to the convexity of the local component function ${f_{i}^{l}}$, $l\in\mathcal{Q}_{i}$. For the sequel analysis, we define respectively the gradient-learning quantities $t_{k}^{u}:=\sum\nolimits_{i\in\mathcal{R}}{t_{i,k}^{u}}$ and $t_{k}^{w}:=\sum\nolimits_{i\in\mathcal{R}}{t_{i,k}^{w}}$ for Prox-DBRO-SAGA and Prox-DBRO-LSVRG, the largest and smallest number of local samples ${q_{\min}}:={\min_{i\in\mathcal{R}}}{q_{i}}$ and ${q_{\max}}:={\max_{i\in\mathcal{R}}}{q_{i}}$, the minimum and maximum triggering probabilities ${p_{\min}}:={\min_{i\in\mathcal{R}}}{p_{i}}$ and ${p_{\max}}:={\max_{i\in\mathcal{R}}}{p_{i}}$, while ${\kappa_{q}}:={q_{\max}}/{q_{\min}}\geq 1$.
@@
(Gradient-learning sequences) Since $t_{k}^{u}$ and $t_{k}^{w}$ are non-negative according to the convexity of the local component function ${f_{i}^{l}}$, $l\in\mathcal{Q}_{i}$ under Assumption 1, $\forall k\geq 0$, we have for Prox-DBRO-SAGA,
@@
where ${D_{F}}\left({{x_{k}},{x^{*}}}\right):=F\left({{x_{k}}}\right)-F\left({{x^{*}}}\right)-\nabla F{\left({{x^{*}}}\right)^{\top}}\left({{x_{k}}-{x^{*}}}\right)$ is known as the Bregman divergence with respect to the convex cost function $F$.
@@
We next seek the upper bound of the distance between the local stochastic gradient estimator $r_{k}$ and gradient $\nabla F\left({{x^{*}}}\right)$ at the optimal solution for both Prox-DBRO-SAGA and Prox-DBRO-LSVRG.
@@
(Gradient-learning errors) Suppose that Assumptions 1-2 hold. For $k\geq 0$, we have for Prox-DBRO-SAGA,
@@
where $r_{k}^{u}:={\mathbf{col}}{\left\{{r_{i,k}^{u}}\right\}_{i\in\mathcal{R}}}\in{\mathbb{R}^{\left|\mathcal{R}\right|n}}$ and $r_{k}^{w}:={\mathbf{col}}{\left\{{r_{i,k}^{w}}\right\}_{i\in\mathcal{R}}}\in{\mathbb{R}^{\left|\mathcal{R}\right|n}}$.
@@
Consider two constants $a_{1}\geq 1$ and $a_{2}>0$, such that $11/{a_{1}}+1/{a_{2}}=$ 1. For an arbitrary vector $\tilde{x}\in{\mathbb{R}^{n}}$, we denote the subdifferential ${\partial}{\left\|{\tilde{x}}\right\|_{a_{1}}}=\left\{{\tilde{z}\in{\mathbb{R}^{n}}:\left\langle{\tilde{z},\tilde{x}}\right\rangle={{\left\|{\tilde{x}}\right\|}_{a_{1}}},{{\left\|{\tilde{z}}\right\|}_{a_{2}}}\leq 1}\right\}$.
@@
where $x={\mathbf{col}}{\left\{{{x_{i}}}\right\}_{i\in\mathcal{R}}}\in{\mathbb{R}^{\left|\mathcal{R}\right|n}}$ and $y={\mathbf{col}}{\left\{{{y_{i}}}\right\}_{i\in\mathcal{R}}}\in{\mathbb{R}^{\left|\mathcal{R}\right|n}}$.
@@ IV-B Main Results
We next derive a feasible range for the penalty parameter to enable the equivalence between the decentralized consensus optimization problem (3) and norm-penalized approximation problem (4) as follows, which further guarantees the equivalence between the original optimization problem (1) and norm-penalized approximation problem (4). We define $\Pi\in{\mathbb{R}^{\left|\mathcal{R}\right|\times\left|{{\mathcal{E}_{\mathcal{R}}}}\right|}}$ by the incidence matrix associated with ${\mathcal{G}_{\mathcal{R}}}$. Specifically, for any edge $e:=\left({i,j}\right)\in{{\mathcal{E}_{\mathcal{R}}}}$ with $i<j$, the $\left({i,e}\right)$-th and $\left({j,e}\right)$-th entries of $\Pi$ are $1$ and $-1$, respectively.
@@
(Resilient consensus condition) Suppose that Assumptions 1 and 2 hold. Given $g^{\prime}\left({{{\tilde{x}}^{*}}}\right)\in{\partial}{g}\left({{{\tilde{x}}^{*}}}\right)$, if the penalty parameter satisfy $\phi\geq{\phi_{\min}}:={\left|\mathcal{R}\right|^{\frac{3}{2}}}\sqrt{\left|{{\mathcal{E}_{\mathcal{R}}}}\right|}\mathop{\max}\nolimits_{i\in\mathcal{R}}{\left\|{\nabla{f_{i}}\left({{{\tilde{x}}^{*}}}\right)}+{g^{\prime}\left({{{\tilde{x}}^{*}}}\right)}\right\|_{\infty}}/{\lambda_{\min}}\left(\Pi\right)$, the optimal solution to the original optimization problem (1) is equivalent to the globally optimal solution to norm-penalized approximation problem (4), i.e., ${x^{*}}={1_{\left|\mathcal{R}\right|}}\otimes{{\tilde{x}}^{*}}$.
@@
Theorem 1 demonstrates that a selection of a sufficiently large penalty parameter guarantees the equivalence between the original optimization problem (1) and norm-penalized approximation problem (4). However, the sequel convergence results manifest that a larger $\phi$ causes a bigger convergence error. Therefore, the notion of a sufficiently large penalty parameter is tailored for theoretical results, and one can hand-tune this parameter to obtain better algorithm performances in practice.
For simplicity, we fix the minimum and maximum heterogenous triggering probabilities as ${p_{\min}}=1/{q_{\max}}$ and $p_{\max}=1/{q_{\min}}$, respectively. Hence, we define ${r_{k}}:=\left\{\begin{gathered}r_{k}^{u},{\text{for {Prox-DBRO-SAGA}}}\hfill\\ r_{k}^{w},{\text{for {Prox-DBRO-LSVRG}}}\hfill\\ \end{gathered}\right.$ and ${t_{k}}:=\left\{\begin{gathered}t_{k}^{u},\text{for {Prox-DBRO-SAGA}}\hfill\\ t_{k}^{w},\text{for {Prox-DBRO-LSVRG}}\hfill\\ \end{gathered}\right.$, such that the theoretical results for both Prox-DBRO-SAGA (Algorithm 2) and Prox-DBRO-LSVRG (Algorithm 3) can be unified in a general framework. Before deriving a linear convergence rate for Algorithms 2-3, we first define the sequel parameters: $\gamma:=\mu L/\left({\mu+L}\right)$, $P_{1}^{c}:=16n{\phi^{2}}\sum\nolimits_{i\in{\mathcal{R}}}{{{\left|{{{\mathcal{R}}_{i}}}\right|}^{2}}}+4n{\phi^{2}}\sum\nolimits_{i\in{\mathcal{R}}}{{{\left|{{{\cal B}_{i}}}\right|}^{2}}}$, $P_{2}:=n{\phi^{2}}\sum\nolimits_{i\in{\mathcal{R}}}{{{\left|{{{\cal B}_{i}}}\right|}^{2}}}/\gamma$, and $E:=4{P_{2}}/\gamma$.
@@
(Linear convergence). Suppose that Assumptions 1-2 hold. Under the conditions of Theorem 1, if the constant step-size meets $0<{\alpha_{k}}\equiv\alpha\leq 1/\left({{\kappa_{q}}\left({32{{\left({1+{\kappa_{f}}}\right)}^{2}}+q_{\text{min}}}\right)\mu}\right)$, then the sequence ${\left\{{{x_{k}}}\right\}_{k\geq 0}}$ generated by Algorithms 2-3, converges linearly to an error ball around the optimal solution to the original optimization problem (1) at a linear rate of ${\left({1-\mathcal{O}\left({\gamma\alpha}\right)}\right)^{k}}$, i.e.,
$\mathbb{E}\left[{\left\|{{x_{k}}-{1_{\left|\mathcal{R}\right|}}\otimes{{\tilde{x}}^{*}}}\right\|_{2}^{2}}\right]$ (13)
@@
We continue to establish the sub-linear convergence rate of Algorithms 2-3 with the aid of the following bounded-gradient assumption, which is standard in literature [50, 51, 52].